# ShEPhERD: diffusing shape, electrostatics, and pharmacophores for bioisosteric drug design

### v1

@@ 3.3 Joint diffusion of molecules and their interaction profiles with ShEPhERD
(1) We collate the coordinates of $\bm{x}_{1}^{(t)},\bm{x}_{2}^{(t)},\bm{x}_{3}^{(t)},$ and $\bm{x}_{4}^{(t)}$ to form a heterogeneous 3D graph where the nodes are attributed with their corresponding latent features $(\bm{z}^{(t)}_{i},\bm{\tilde{z}}^{(t)}_{i})\ \forall\ i\in[1,2,3,4]$. We then encode this heterogeneous 3D graph with another E3NN (EquiformerV2) module $\phi_{joint}^{local}$ and residually update the nodes’ latent features: $(\bm{z}^{(t)}_{i},\bm{\tilde{z}}^{(t)}_{i})\mathrel{{+}{=}}\phi_{joint}^{local}\bigg{(}(\bm{x}_{i}^{(t)},\bm{z}^{(t)}_{i},\bm{\tilde{z}}^{(t)}_{i})\ \forall\ i\in[1,2,3,4]\bigg{)}$.
(2) We sum-pool the updated $l{=}1$ node features across each sub-graph, concatenate, and then embed with an equivariant feed-forward network $\phi_{joint}^{global}$ to obtain a global $l{=}1$ code describing the overall system: $\bm{\tilde{z}}_{joint}^{(t)}=\phi_{joint}^{global}\bigg{(}\text{Cat}\bigg{[}\bigg{(}\sum_{k=1}^{n_{i}}\bm{\tilde{z}}^{(t)}_{i}[k]\bigg{)}\ \forall\ i\in[1,2,3,4]\bigg{]}\bigg{)}$. We then apply equivariant tensor products between $\bm{\tilde{z}}_{joint}^{(t)}\in\mathbb{R}^{d\times 3}$ and an $l{=}0$ embedding of $t$. This yields $l{=}0$ and $l{=}1$ global latent features $(\bm{z}_{joint}^{(t)},\bm{\tilde{z}}_{joint}^{(t)})$, which are residually added to the node representations $(\bm{z}^{(t)}_{i},\bm{\tilde{z}}^{(t)}_{i})$.
@@
Sampling. For unconditional generation, we first sample $\bm{X}^{(T)}$ from isotropic Gaussian noise, and then denoise for $T$ steps to sample $(\bm{x}_{1}^{(0)},\bm{x}_{2}^{(0)},\bm{x}_{3}^{(0)},\bm{x}_{4}^{(0)})$. We then argmax $\bm{a}^{(0)}$, $\bm{f}^{(0)}$, $\bm{B}^{(0)}$, and $\bm{p}^{(0)}$ to obtain discrete atom/bond/pharm. types, and round each $\bm{V}^{(0)}[k]$ to have norm 1.0 or 0.0. To help break the spherical symmetry of the coordinates of $\bm{X}^{(t)}$ at early time steps, we find it useful to strategically add extra noise to the reverse process (App. A.8). For conditional generation, we use inpainting to sample $\bm{x}_{1}^{(0)}$ conditioned on target interaction profiles. Namely, we first simulate the forward-noising of the target profiles $(\bm{x}^{*}_{2},\bm{x}^{*}_{3},\bm{x}^{*}_{4})\rightarrow(\bm{x}^{*}_{2},\bm{x}^{*}_{3},\bm{x}^{*}_{4})^{(t)}\ \forall\ t\in[1,T]$. Then, during the reverse denoising process, we replace the ShEPhERD-denoised $(\bm{x}_{2}^{(t)},\bm{x}_{3}^{(t)},\bm{x}_{4}^{(t)})$ with the noisy target $(\bm{x}^{*}_{2},\bm{x}^{*}_{3},\bm{x}^{*}_{4})^{(t)}$. Like other molecule DDPMs, we must specify $n_{1}$ and $n_{4}$ for each sample.
@@ 4 Experiments
Bioactive hit diversification. Whereas stucture-based drug design aims to design high-affinity ligands given the structure of the protein target, ligand-based drug design attempts to develop and optimize bioactive hit compounds in the absence of protein information. A common task in ligand-based drug design is diversifying the chemical structures of previously identified bioactive hits (i.e., from a phenotypic experimental screen) through interaction-preserving scaffold hopping, often as a means to reduce toxicity, increase synthesizability, or evade patent restrictions. We simulate bioactive hit diversification by using ShEPhERD to generate analogues of 7 experimental ligands from the PDB. To evaluate their likelihood of retaining bioactivity, we use Autodock Vina (Trott & Olson, 2010; Eberhardt et al., 2021) to dock the generated ligands to their respective proteins, treating the Vina docking scores as a weak surrogate for bioactivity. To best imitate ligand-based design, we condition ShEPhERD on the ESP surface and pharmacophores of the lowest-energy conformer of each PDB ligand, rather than their bound poses (we also simulate this scenario for comparison). We then use inpainting to generate 500 samples from $P(\bm{x}_{1}|\bm{x}_{3},\bm{x}_{4})$, and dock the valid samples. Fig. 4 shows the distributions of Vina scores for ShEPhERD-generated analogues, compared against a docking screen of 10K random compounds from ShEPhERD-MOSES-aq. Despite having no knowledge of the protein targets, ShEPhERD enriches average and top-10 Vina scores in 5/7 cases. For the top-scoring generated ligands for 5mo4 and 7l11, ShEPhERD generates substantial scaffold hops that lead to diverse 2D graph structures relative to the experimental ligands. Nevertheless, upon docking, the generated molecules still explore poses that closely align with the experimental crystal poses.
@@ Acknowledgments
This research was supported by the Office of Naval Research under grant number ONR N00014-21-1-2195. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. 2141064. The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing HPC resources that have contributed to the research results reported within this paper.
@@ A.3 Defining and analyzing validity of generated 3D molecules
Table 3 and Figures 11-12 show the strain energy and heavy-atom RMSD of ShEPhERD-generated molecular structures compared to their xTB-relaxed geometries. Note that RMSDs are computed after realigning (by minimizing RMSD) the relaxed geometry to the unrelaxed geometry. Mean strain energies remain relatively low for the model trained on ShEPhERD-GDB17 (mean strain energies ${<}10$ kcal/mol). Indeed, these strain energies are typically lower than the strain energies of molecules with MMFF94-optimized geometries that have their heavy-atom RMSD and strain energies measured after relaxing the MMFF94 geometries with xTB. However, the high standard deviations in the strain energy distributions indicate that there are still many generated molecules that are highly strained (up to nearly 1000 kcal/mol). Nevertheless, almost all of the RMSD’s are ${<}1$Å which implies that the generated molecules are close to the true xTB geometry. Indeed, the unconditionally-generated samples for models trained on ShPhERD-GDB17 have average heavy-atom RMSDs below $0.08$. We note that slightly distorted atomic positions can vastly increase strain energy, which makes strain energy an unforgiving (but still useful) measure of absolute geometric quality. We notice a similar trend to validity where samples from $P(\bm{x}_{1},\bm{x}_{2})$ model perform the best in terms of strain energy while samples from $P(\bm{x}_{1},\bm{x}_{4})$ model perform the worst, though the RMSD’s are quite low across all the unconditional models.
@@ A.10 Additional examples of generated molecules
Figures 22, 23, and 24 show additional random examples of unconditionally-generated molecules and their jointly generated interaction profiles from versions of ShEPhERD trained on ShEPhERD-GDB17. Figures 25, 26, and 27 show additional random examples of ShEPhERD-generated molecules when conditioning on target interaction profiles via inpainting. Although we are primarily interested in using the models trained to learn $P(\bm{x}_{1},\bm{x}_{3},\bm{x}_{4})$ on ShEPhERD-MOSES-aq for conditional generation tasks relevant to ligand-based drug design, we also show unconditional samples from this model in Fig. 28.

### v2

@@ 3.3 Joint diffusion of molecules and their interaction profiles with ShEPhERD
(1) We collate the coordinates of $\bm{x}_{1}^{(t)},\bm{x}_{2}^{(t)},\bm{x}_{3}^{(t)},$ and $\bm{x}_{4}^{(t)}$ to form a heterogeneous 3D graph where the nodes are attributed with their corresponding latent features $(\bm{z}^{(t)}_{i},\bm{\tilde{z}}^{(t)}_{i})\ \forall\ i\in[1,2,3,4]$. We then encode this heterogeneous 3D graph with another E3NN (EquiformerV2) module $\phi_{joint}^{local}$ and residually update the nodes’ latent features: $(\bm{z}^{(t)}_{i},\bm{\tilde{z}}^{(t)}_{i})\mathrel{{+}{=}}\phi_{joint}^{local}\left((\bm{x}_{i}^{(t)},\bm{z}^{(t)}_{i},\bm{\tilde{z}}^{(t)}_{i})\ \forall\ i\in[1,2,3,4]\right)$.
(2) We sum-pool the updated $l{=}1$ node features across each sub-graph, concatenate, and then embed with an equivariant feed-forward network $\phi_{joint}^{global}$ to obtain a global $l{=}1$ code describing the overall system: $\bm{\tilde{z}}_{joint}^{(t)}=\phi_{joint}^{global}\left(\text{Cat}\left[\left(\sum_{k=1}^{n_{i}}\bm{\tilde{z}}^{(t)}_{i}[k]\right)\ \forall\ i\in[1,2,3,4]\right]\right)$. We then apply equivariant tensor products between $\bm{\tilde{z}}_{joint}^{(t)}\in\mathbb{R}^{d\times 3}$ and an $l{=}0$ embedding of $t$. This yields $l{=}0$ and $l{=}1$ global latent features $(\bm{z}_{joint}^{(t)},\bm{\tilde{z}}_{joint}^{(t)})$, which are residually added to the node representations $(\bm{z}^{(t)}_{i},\bm{\tilde{z}}^{(t)}_{i})$.
@@
Sampling. For unconditional generation, we first sample $\bm{X}^{(T)}$ from isotropic Gaussian noise, and then denoise for $T$ steps to sample $\bm{X}^{(0)}$. We then argmax $\bm{a}^{(0)}$, $\bm{f}^{(0)}$, $\bm{B}^{(0)}$, and $\bm{p}^{(0)}$ to obtain discrete atom/bond/pharm. types, and round each $\bm{V}^{(0)}[k]$ to have norm 1.0 or 0.0. To help break the spherical symmetry of $\bm{X}^{(t)}$ at early time steps, we strategically add extra noise to the reverse process (App. A.8). For conditional generation, we use inpainting (Lugmayr et al., 2022; Schneuing et al., 2022) to sample $\bm{x}_{1}^{(0)}$ conditioned on target interaction profiles. Namely, we first simulate the forward-noising of the target profiles $(\bm{x}^{*}_{2},\bm{x}^{*}_{3},\bm{x}^{*}_{4})\rightarrow(\bm{x}^{*}_{2},\bm{x}^{*}_{3},\bm{x}^{*}_{4})^{(t)}\ \forall\ t\in[1,T]$. Then, during the reverse denoising process, we replace the ShEPhERD-denoised $(\bm{x}_{2}^{(t)},\bm{x}_{3}^{(t)},\bm{x}_{4}^{(t)})$ with the noisy target $(\bm{x}^{*}_{2},\bm{x}^{*}_{3},\bm{x}^{*}_{4})^{(t)}$. Like other molecule DDPMs, we must specify $n_{1}$ and $n_{4}$ for each sample.
@@ 4 Experiments
Bioactive hit diversification. Whereas stucture-based drug design aims to design high-affinity ligands given the structure of the protein target, ligand-based drug design attempts to develop and optimize bioactive hit compounds in the absence of protein information. A common task in ligand-based drug design is diversifying the chemical structures of previously identified bioactive hits (i.e., from a phenotypic experimental screen) through interaction-preserving scaffold hopping, often as a means to reduce toxicity, increase synthesizability, or evade patent restrictions. We simulate bioactive hit diversification by using ShEPhERD to generate analogues of 7 experimental ligands from the PDB. To evaluate their likelihood of retaining bioactivity, we use Autodock Vina (Trott & Olson, 2010; Eberhardt et al., 2021) to dock the generated ligands to their respective proteins, treating the Vina docking scores as a weak surrogate for bioactivity. To best imitate ligand-based design, we condition ShEPhERD on the ESP surface and pharmacophores of the lowest-energy conformer of each PDB ligand, rather than their bound poses (we also simulate this scenario for comparison). We then use inpainting to generate 500 samples from $P(\bm{x}_{1}|\bm{x}_{3},\bm{x}_{4})$, and dock the valid samples. Fig. 4 shows the distributions of Vina scores for ShEPhERD-generated analogues, compared against a docking screen of 10K random compounds from ShEPhERD-MOSES-aq. Despite having no explicit knowledge of the protein targets, ShEPhERD enriches Vina scores in multiple cases. For the top-scoring generated ligands for 5mo4 and 7l11, ShEPhERD generates substantial scaffold hops that yield diverse 2D graph structures relative to the experimental ligands. Nevertheless, upon docking, the generated molecules still explore poses that closely align with the experimental crystal poses.
@@ Acknowledgments
The authors would like to thank Wenhao Gao for support with SynFormer. This research was supported by the Office of Naval Research under grant number ONR N00014-21-1-2195. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. 2141064. The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing HPC resources that have contributed to the research results reported within this paper.
@@ A.1 Comparing ShEPhERD to REINVENT and virtual screening for natural product ligand hopping
A.1.1 REINVENT convergence
We plot the average score of each REINVENT batch vs. the iteration number for each natural product target in Fig. 6. For the first natural product, REINVENT appears converged. For the second, REINVENT appears close to converged. For the third natural product, REINVENT could likely find better scoring molecules if trained for longer. We emphasize that we could also run ShEPhERD for longer, too. Although REINVENT may not necessarily be converged for each natural product, we follow the PMO benchmark (Gao et al., 2022) and limit the number of oracle calls to 10,000. In practice, even allowing for just 10,000 REINVENT samples required multiple days of computation per target because of the expensive conformer sampling, xTB optimization, and alignment with our scoring functions. We also note that in contrast to REINVENT, ShEPhERD is not specifically trained to optimize similarity for any natural product (or for any particular molecule). Fig. 6 also plots SA score distributions for both REINVENT- and ShEPhERD-generated molecules.
Figure 6: (Top) The average score (ESP+pharmacophore 3D similarity) across a batch for each iteration of the optimization for REINVENT. Only natural product (1) seems to have fully converged. All optimizations were concluded at 10,000 oracle calls. (Bottom) The distributions of SA score for REINVENT and ShEPhERD for molecules with SA score $<4.5$.
@@ A.2.4 Bioactive hit diversification
Figure 10: The individually visualized docking poses for each sampled molecule for PDB targets 5mo4 and 7l11 from Fig. 4.
@@ A.3 Defining and analyzing validity of generated 3D molecules
Table 3 and Figures 13-14 show the strain energy and heavy-atom RMSD of ShEPhERD-generated molecular structures compared to their xTB-relaxed geometries. Note that RMSDs are computed after realigning (by minimizing RMSD) the relaxed geometry to the unrelaxed geometry. Mean strain energies remain relatively low for the model trained on ShEPhERD-GDB17 (mean strain energies ${<}10$ kcal/mol). Indeed, these strain energies are typically lower than the strain energies of molecules with MMFF94-optimized geometries that have their heavy-atom RMSD and strain energies measured after relaxing the MMFF94 geometries with xTB. However, the high standard deviations in the strain energy distributions indicate that there are still many generated molecules that are highly strained (up to nearly 1000 kcal/mol). Nevertheless, almost all of the RMSD’s are ${<}1$Å which implies that the generated molecules are close to the true xTB geometry. Indeed, the unconditionally-generated samples for models trained on ShPhERD-GDB17 have average heavy-atom RMSDs below $0.08$. We note that slightly distorted atomic positions can vastly increase strain energy, which makes strain energy an unforgiving (but still useful) measure of absolute geometric quality. We notice a similar trend to validity where samples from $P(\bm{x}_{1},\bm{x}_{2})$ model perform the best in terms of strain energy while samples from $P(\bm{x}_{1},\bm{x}_{4})$ model perform the worst, though the RMSD’s are quite low across all the unconditional models.
@@ A.10 Additional examples of generated molecules
Figures 24, 25, and 26 show additional random examples of unconditionally-generated molecules and their jointly generated interaction profiles from versions of ShEPhERD trained on ShEPhERD-GDB17. Figures 27, 28, and 29 show additional random examples of ShEPhERD-generated molecules when conditioning on target interaction profiles via inpainting. Although we are primarily interested in using the models trained to learn $P(\bm{x}_{1},\bm{x}_{3},\bm{x}_{4})$ on ShEPhERD-MOSES-aq for conditional generation tasks relevant to ligand-based drug design, we also show unconditional samples from this model in Fig. 30.
@@ A.11 Training and inference resources
A.12 Additional Experiments and Comparisons to Related Work
A.12.1 Comparison to SQUID for shape-conditioned generation
We compare ShEPhERD to two ligand-based models, SQUID (Adams & Coley, 2022) and ShapeMol (Chen et al., 2023), for shape-conditioned molecular design. To perform this comparison, we apply the version of ShEPhERD trained on MOSES-aq to learn $P(\bm{x}_{1},\bm{x}_{3},\bm{x}_{4})$ to generate new molecules conditioned on the surface shape (e.g., the coordinates of $\bm{x}_{3}$, which implicitly define $\bm{x}_{2}$) of the 1000 3D molecules in SQUID’s test set. Note that SQUID and ShEPhERD use the same training and test splits of MOSES, and hence there was no data leakage. We condition the $P(\bm{x}_{1},\bm{x}_{3},\bm{x}_{4})$ model on the molecules’ shapes (only) by inpainting just the coordinates of $\bm{x}_{3}$, and allowing ShEPhERD to generate $\bm{x}_{1}$, $\bm{x}_{4}$, and the electrostatic potentials of $\bm{x}_{3}$. We generate up to 10 samples from ShEPhERD and compare the average volumetric shape similarity using ROCS (the same shape scoring function that SQUID used). Since SQUID’s generated molecules for each test molecule are publicly available for download via their GitHub, we re-evaluated the average shape similarity of SQUID-generated molecules to ensure a fair comparison.
Across the 1000 test-set molecules, the average shape similarity of SQUID-generated molecules is $0.70$ when using $\lambda=1.0$, and $0.74$ when using $\lambda=0.3$. In contrast, the average shape similarity of ShEPhERD-generated molecules is $0.80$, a substantial improvement. Moreover, the ShEPhERD-generated molecules have an average 2D graph similarity of just 0.22 compared to the reference molecule, whereas the SQUID-generated molecules have average graph similarities of $0.25$ ($\lambda=1.0$) and $0.35$ ($\lambda=0.3$). Hence, ShEPhERD is able to generate less chemically similar molecules that have higher shape similarity to the target, on average. The full results are found in Table 10.
We emphasize that ShEPhERD greatly exceeds the performance of SQUID even with this indirect way of doing shape-conditioned generation, i.e. via inpainting the joint $P(\bm{x}_{1},\bm{x}_{3},\bm{x}_{4})$ model. We could also train a $P(\bm{x}_{1},\bm{x}_{2})$ model directly for this task; however, we did not have time to retrain ShEPhERD from scratch within the rebuttal time period. We also note that ShEPhERD conditions on surface shapes, whereas we have compared ShEPhERD to SQUID using volumetric shape similarity (with ROCS). We expect that we could further improve ShEPhERD’s performance in shape-conditioned generation by retraining ShEPhERD with a smaller surface probe radius, thereby “shrinking” the surface shape to better model volumetric shapes.
Table 10: A comparison of ShEPhERD to default versions of SQUID ($\lambda=\{1.0,0.3\}$) and ShapeMol. We recompute the shape and graph similarities of SQUID using the structures provided by their GitHub. Note that the values for the evaluations of ShapeMol and ShapeMol+g were directly copied from (Chen et al., 2023) which use a different alignment protocol.
Model $\text{{avgSim}}_{s}$ ($\uparrow$) $\text{{avgSim}}_{g}$ ($\downarrow$) QED ($\uparrow$)
ShEPhERD $\mathbf{0.799}\pm 0.058$ $\mathbf{0.223}\pm 0.065$ $0.723$
SQUID ($\lambda=1.0$) $0.700\pm 0.107$ $0.245\pm 0.078$ $0.760$
SQUID ($\lambda=0.3$) $0.735\pm 0.117$ $0.346\pm 0.161$ $\mathbf{0.766}$
ShapeMol $0.689\pm 0.044$ $0.239\pm 0.042$ $0.748$
ShapeMol+g $0.746\pm 0.036$ $0.241\pm 0.050$ $0.749$
A.12.2 Comparisons against structure-based drug design models that explicitly encode the protein pocket
While ShEPhERD is a ligand-only model and does not explicitly model protein pocket geometries, we were asked to compare to structure-based drug design (SBDD) models. We emphasize that since the methods differ in their respective applications and modeling strategies, the comparisons cannot be completely fair. Nevertheless, although ShEPhERD is not designed for structure-based drug design, it can still be roughly applied to SBDD by conditioning on the bound pose of a known ligand (but still ignoring the surrounding protein pocket). This bound pose can be obtained from crystallography (e.g., the co-crystal ligands in PDB entries) or from simulated docked poses (e.g., conditioning ShEPhERD on the top hit from a small docking-based virtual screen). In order to compare to SBDD generative models, we compare ShEPhERD against the benchmarking results presented by (Zheng et al., 2024), which report the average Vina docking scores amongst the top-10 generated compounds (amongst $\sim$1000 generated samples) for 7 PDB systems using generative SBDD methods like 3DSBDD, AutoGrow4, Pocket2Mol, PocketFlow, and ResGen. Specifically, we use ShEPhERD to condition on the electrostatic potential surface ($x_{3}$) and pharmacophores ($x_{4}$) of the PDB ligand in its crystallographic bound pose, generating up to 1000 valid molecules per PDB target. We dock the 1000 generated molecules (starting from their SMILES strings) using the same Vina program (and version) as used by Zheng et al. (2024), and compute the average Vina scores amongst the top-10 molecules per target. These average top-10 scores are compared against the results for the SBDD models evaluated by Zheng et al. (2024) in Table 11. We also evaluate the 2D graph similarities of ShEPhERD-generated compounds compared to the PDB ligand in each case, which are generally quite low (Fig. 31).
Table 11: Docking scores of ShEPhERD-generaed molecules and redocked PDB ligands, compared to ligands generated by various SBDD generative models. Scores for 3DSBDD, AutoGrow, Pocket2Mol, PocketFlow, and ResGen are copied from Zheng et al. (2024).
Model 1iep 3eml 3ny8 4rlu 4unn 5mo4 7l11
Redocked PDB ligand -9.58 -9.14 -8.62 -8.70 -9.24 -9.95 -9.10
ShEPhERD -12.25 -10.80 -10.69 -10.19 -10.60 -10.66 -9.44
± 0.18 ± 0.23 ± 0.23 ± 0.23 ± 0.27 ± 0.20 ± 0.19
3DSBDD -9.05 -10.02 -10.10 -9.80 -8.23 -8.71 -8.47
± 0.38 ± 0.15 ± 0.24 ± 0.55 ± 0.30 ± 0.45 ± 0.18
AutoGrow4 -13.23 -13.03 -11.70 -11.20 -11.14 -10.38 -8.84
± 0.11 ± 0.09 ± 0.00 ± 0.00 ± 0.12 ± 0.27 ± 0.33
Pocket2Mol -10.17 -12.25 -11.89 -10.57 -12.20 -10.07 -9.74
± 0.53 ± 0.27 ± 0.16 ± 0.12 ± 0.34 ± 0.62 ± 0.38
PocketFlow -12.49 -9.25 -8.56 -9.65 -7.90 -7.80 -8.35
± 0.70 ± 0.29 ± 0.35 ± 0.25 ± 0.78 ± 0.42 ± 0.31
ResGen -10.97 -9.25 -10.96 -11.75 -9.41 -10.34 -8.74
± 0.29 ± 0.95 ± 0.42 ± 0.42 ± 0.23 ± 0.39 ± 0.24
Even though ShEPhERD only sees the PDB ligand and not the target protein pocket, the average top-10 Vina scores for ShEPhERD-generated molecules are very comparable to those obtained by these SBDD methods. Out of the 6 generative models (ShEPhERD and 5 SBDD methods), ShEPhERD ranks 1/6 for one target, 2/6 for one target, 3/6 for three targets, and 4/6 for two targets. These results confirm that ShEPhERD shows great promise for SBDD even though it is currently a ligand-only model.
Figure 31: The graph similarity distributions of the 1000 valid ShEPhERD-generated molecules used in App. A.12.2 for each target. We also show the distributions of 1000 randomly selected samples from the 10k docking screen. Note that only 672 ShEPhERD-generated molecules were valid (and docked) for the target 1iep.
Figure 32: The graph similarity distributions of the valid ShEPhERD-generated molecules used in the main text. We also show the distributions of 500 randomly selected samples from the 10k docking screen.
A.12.3 Comparisons against inpainting with DiffSBDD
We also directly compare ShEPhERD to DiffSBDD, which also uses diffusion with inpainting to generate molecules, albeit in a structure-based drug design scenario where the pocket is directly encoded by the model. To compare ShEPhERD specifically against DiffSBDD, we applied ShEPhERD to generate analogs given the crystallographic bound pose for the ligand in the PDB entry 4tos (ligand PDB ID: 355), which DiffSBDD was evaluated on in Figure 2 of Schneuing et al. (2022). To ensure a fair comparison, we downloaded the 100 molecules generated by DiffSBDD for this PDB target, which are fortunately made publicly available by Schneuing et al. (2022). We used the molecules generated by the all-atom version of DiffSBDD that uses inpainting, as this is the model that generally performs the best according to Schneuing et al. (2022). We then re-docked these 100 molecules (starting from their SMILES strings) with AutoDock Vina (v1.1.2) to obtain their distribution of docking scores. We then applied ShEPhERD to generate 100 valid molecules given ($\mathbf{x}_{3}$, $\mathbf{x}_{4}$) of the bound ligand (still ignoring the actual protein pocket), and docked those molecules, again starting from their SMILES strings.
We compare the distributions of docking scores between the molecules generated by ShEPhERD and DiffSBDD in Figure 33. Even though ShEPhERD does not explicitly model the pocket structure, the average Vina score for ShEPhERD-generated molecules is approximately 2 kcal/mol lower (better) than the average Vina score for the DiffSBDD-generated molecules. Moreover, the ShEPhERD-generated molecules have significantly lower (better) SA scores, as well. Meanwhile, the ShEPhERD-generated molecules still have quite low graph similarity to the PDB ligand (vast majority are below 0.2). We also emphasize that this conditioning PDB ligand is clearly out-of-distribution for ShEPhERD, as the PDB ligand contains 42 non-hydrogen atoms, whereas ShEPhERD is only trained on molecules containing up to 27 non-hydrogen atoms. However, we do want to emphasize that, as previously mentioned, comparisons to SBDD methods are inherently unfair. In this case, the PDB ligand itself has a quite good docking score, which means that the ShEPhERD-generated analogs are likely to also have good docking scores, as ShEPhERD generates molecules that preserve 3D interactions. Nevertheless, this experiment shows ShEPhERD’s utility in SBDD, despite being a ligand-only model.
Figure 33: A comparison of distributions for molecules generated by DiffSBD and ShEPhERD for the target 4tos (PDB ligand ID 355). We show distributions of Vina Score (v1.1.2), SA score, and 2D graph similarity. (Top) The distributions of all 100 valid molecules generated by both models. (Bottom) The distributions of the top-10 docked molecules from both models.
A.12.4 Comparisons against SynFormer on natural product ligand hopping
We compare against SynFormer (Gao et al., 2024) in our natural product ligand hopping experiments. However, note that the goal of our natural product ligand hopping experiments was to generate small-molecule analogues of structurally complex natural products that (1) preserve their electrostatic and pharmacophore 3D interactions, and (2) have simpler chemical structures as assessed via SA score. SynFormer is a non-3D generative model that can be used to generate synthesizable small-molecule analogues of an unsynthesizable reference molecule that have similar 2D chemical structures. We applied SynFormer using its publicly available code in its default settings to generate analogs for each of the three natural products. We took each of the generated analogs, generated up to 10 xTB-optimized conformers, and scored their 3D similarities to the reference natural products after optimal alignment.
We compare the SynFormer-generated analogs against the ShEPhERD-generated analogs in Figure 34 on the basis of electrostatic and pharmacophore similarity to the reference natural products. In all cases, SynFormer underperforms ShEPhERD, which is not surprising since SynFormer was not designed to preserve 3D interaction similarity. We also note that very few (if any) of the SynFormer-generated molecules have SA scores below the threshold we enforce on ShEPhERD (SA$<$4.5), indicating that SynFormer’s analogs are generally more complex, even if purportedly synthesizable.
Figure 34: (Top) Plots of ESP vs. pharmacophore similarity for the molecules generated by SynFormer and ShEPhERD in our natural product ligand hopping tasks. (Bottom) SA score distributions for each model’s generated molecules.